OpenStudy (anonymous):

Is (5, 17) a solution of the equation y = 4x + 3? A. yes B. no

jimthompson5910 (jim_thompson5910):

plug in x = 5 and y = 17 and tell me what you get

jimthompson5910 (jim_thompson5910):

do you know what I mean?

OpenStudy (anonymous):

Let y=17 and let x=5 17=4*5+3 17=20+3 17=23 <- this is not true. The answer is No.